### v4

# No Other Representation Component Is Needed: Diffusion Transformers Can Provide Representation Guidance by Themselves

Abstract
Recent studies have demonstrated that learning a meaningful internal representation can both accelerate generative training and enhance the generation quality of diffusion transformers. However, existing approaches necessitate to either introduce an external and complex representation training framework or rely on a large-scale, pre-trained representation model to provide representation guidance during the original generative training process. In this study, we posit that the unique discriminative process inherent to diffusion transformers enables them to offer such guidance without requiring external representation components. We therefore propose Self-Representation Alignment (SRA), a simple yet effective method that obtains representation guidance through a self-distillation manner. Specifically, SRA aligns the latent representation of the diffusion transformer in the earlier layer conditioned on higher noise to that in the later layer conditioned on lower noise to progressively enhance the overall representation learning during only the generative training process. Experimental results indicate that applying SRA to DiTs and SiTs yields consistent performance improvements. Moreover, SRA not only largely outperforms approaches relying on auxiliary, complex representation training frameworks but also achieves performance comparable to methods that are heavily dependent on an external pre-trained representation model. Refer to caption Figure 1: Left: Methods like MaskDiT [84] and SD-DiT [86] use an external representation task to guide diffusion transformer. Middle: Methods like REPA [81] leverage an external representation foundation model as guidance. Right (our approach): We do not use any external representation component but still obtain such guidance through proposed self-representation alignment technique.
1 Introduction
Diffusion transformers [62, 56, 10] and vision transformers [20, 52, 75] have held the dominant positions in visual generation and representation because of their scalability during pre-training [5, 22, 61, 72] and generalization capacity for downstream tasks [41, 50, 38, 48].
Recently, many works [84, 86, 81, 42] have explored leveraging representation components of vision transformers for diffusion transformer’s training and have shown that learning a high-quality internal representation can not only speed up the generative training progress but also improve the generation quality. These works either utilize the training paradigm in representation learning (e.g., MAE’s [29], IBOT’s [85]) shown in Figure 1a or leverage a large-scale pre-trained representation foundation model (e.g., DINOv2 [61], CLIP [65]) shown in Figure 1b to give representation guidance for the diffusion transformer during the original generative training. However, the former method requires a complex training framework design, and the latter method relies on a powerful model, which is trained on a massive amount of data with hundreds of GPUs. Thereout, an intriguing yet underexplored problem has come to light: Can we obtain such representation guidance only during the generative training without external representation components?
Figure 2: We empirically investigate the representations in diffusion transformers across different blocks and timesteps with the original SiT-XL/2 checkpoint trained for 7M iterations. Left: Using PCA [1] to visualize the latent features in SiT, we observe that the features lead a process from coarse to fine when increasing block layers and decreasing noise level. Right: A similar trend can also be seen in the linear probing results on ImageNet. Investigation of DiT is provided in Appendix A, which leads to the similar results as SiT.
Our observations: Different from the representation model that takes a clean image as input and then outputs semantically-rich feature, diffusion model often takes a noise latent as input and obtains cleaner one step-by-step. In other words, the generative mechanism by which the diffusion model operates can be generally considered as a coarse to fine process. Inspired by this behavior, we hypothesise that the representations in it also follow such a trend. To testify this, we perform an empirical analysis with recent diffusion transformers [56, 62]. As shown in Figure 2a, we first find out that the latent features in the diffusion transformer are progressively refined, moving from coarse to fine, as block layers increase and noise level decreases. Next, akin to the results in previous studies [81, 77], we observe that the diffusion transformer already learns meaningful discriminative representations as shown in Figure 2b. Meanwhile, although the accuracy drops off after reaching a peak at about layer 20 because the model needs to shift away to focus on generating images with high-frequency details, the quality of the representations basically transfers from bad to good by increasing block layers and decreasing noise level. These indicate that the diffusion transformer gets roughly from a coarse-to-fine discriminative process when only generative training is performed.
This distinctive trend motivates us to align the weaker representations in the diffusion transformer to the better ones in the original generative training, thereby enhancing the representation learning of the model without involving any external representation component. However, this approach is not straightforward during training. As observed in previous studies [12, 28], directly using the output of the same model as a supervision signal may lead to shortcut learning [26], which can compromise performance or even cause training collapse.
Our approach: To overcome the abovementioned hurdles, we present Self-Representation Alignment (SRA), a simple but effective technique built on recent diffusion transformer architectures [56, 62]. As shown in Figure 1c, SRA does not need any representation component; in essence, it aligns the output latent representation in earlier layer conditioned on higher noise to that in later layer conditioned on lower noise to achieve self-representation alignment. Meanwhile, in order to make the training process more stable, we obtain the target features from another model that shares the same architecture with the trainable model but updates weight by weighted moving average (EMA)***In the following part, we use term ‘student’ to represent the trainable model and ‘teacher’ to represent the EMA model in SRA for simplicity.. Furthermore, the student’s output latent feature is first passed through the projection layers to conduct a slight nonlinear transformation for better representation extraction and then aligned with the target feature output by the teacher. In a nutshell, our SRA can offer a flexible way to integrate representation guidance without external component needs and architect modification.
Finally, we conduct comprehensive experiments to evaluate the effect of SRA. After a series of component-wise analyses within the design space, we show that SRA brings significant performance improvements to both DiTs [62] and SiTs [56]. Moreover, our ablation study highlights the crucial role of internal representations in the success of SRA, which supports our central hypothesis: diffusion transformers can achieve representation alignment without external components.
In summary, our main contributions are as follows:
  • We analyze the latent representation in the diffusion transformers and assume that the unique coarse to fine discriminative process makes it possible to achieve representation alignment without external components.
  • We introduce SRA, a simple yet effective method that aligns the output latent representation of the diffusion transformers in the earlier layer conditioned on higher noise to that in the later layer conditioned on lower noise to achieve self-representation alignment.
  • With our SRA, both DiTs and SiTs achieve sustained training speed acceleration and nontrivial generation performance improvement.
2 Method
2.1 Preliminary: Training Object of DiT and SiT
As our method is built upon the denoise-based model (DiT) and flow-based model (SiT), to facilitate a more seamless introduction of SRA in the subsequent part, we now present a brief overview of the training object of these two types of models. We omit the class-condition here for simplicity because it can be easily added to all analogs without causing any changes to the mathematical equations. We leave more detailed mathematical descriptions of these types of models in Appendix B.
Denoise-based models learn to transform Gaussian noise into data samples through a step-by-step denoising process. Given a pre-defined forward process that gradually adds noise, these models learn the reverse process to recover the original data.
For data point ${\mathbf{x}}_{0}$ from distribution ${\mathbf{x}}_{0}\sim p({\mathbf{x}})$, the forward process follows: $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})=\mathcal{N}({\mathbf{x}}_{t};\sqrt{1-\beta_{t}}{\mathbf{x}}_{0},\beta_{t}\mathbf{I})$. The model learns to reverse this process using a neural network $\bm{\epsilon}_{\bm{\zeta}}({\mathbf{x}}_{t},t)$ that predicts the noise added at each step. The network is trained using a simple mean squared error objective that measures how well it can predict the noise:
$\mathcal{L}_{\text{simple}}=\mathbb{E}{{\mathbf{x}}_{t},\bm{\epsilon},t}\Big{[}||\bm{\epsilon}-\bm{\epsilon}_{\bm{\zeta}}({\mathbf{x}}_{t},t)||_{2}^{2}\Big{]}.$ (1)
Different from the denoise-based model, the flow-based model aims to learn a velocity field ${\mathbf{v}}_{\zeta}({\mathbf{x}}_{t},t)$ that governs a probability flow ordinary differential equation (PF ODE). This PF ODE allows the model to sample data by flowing toward the data distribution. This forward process is defined as:
${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{0}+\sigma_{t}{\mathbf{\epsilon}},\quad\alpha_{0}=\sigma_{T}=1,\alpha_{T}=\sigma_{0}=0,$ (2)
where ${\mathbf{x}}_{0}\sim p({\mathbf{x}})$ is the data, ${\mathbf{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is Gaussian noise, and $\alpha_{t}$ and $\sigma_{t}$ are monotone decreasing and increasing functions of $t\in[0,T]$, respectively. The PF ODE is given by:
$\dot{{\mathbf{x}}}_{t}={\mathbf{v}}_{\bm{\zeta}}({\mathbf{x}}_{t},t),$ (3)
Where the marginal distribution of this ODE at time $p_{t}({\mathbf{x}})$ matches the marginal distribution of the forward process. To learn the velocity field, the model ${\mathbf{v}}_{\zeta}({\mathbf{x}}_{t},t)$ is trained to minimize the following loss function:
$\mathcal{L}_{\text{velocity}}=\mathbb{E}_{{\mathbf{x}}_{0},\bm{\epsilon},t}\big{[}||{\mathbf{v}}_{\bm{\zeta}}({\mathbf{x}}_{t},t)-\dot{\alpha}_{t}{\mathbf{x}}_{0}-\dot{\sigma}_{t}\bm{{\mathbf{\epsilon}}}||^{2}\big{]}.$ (4)
For the sake of simplicity, in the following part, we use generative loss ($\mathcal{L}_{\text{gen}}$) to uniformly represent the two generative training objectives of denoise-based and flow-based methods.
2.2 Self-Representation Alignment
Figure 3: Overall framework. SRA aligns the student’s latent representation in the earlier layer conditioned on higher noise (green branch) to that of the teacher in the later layer conditioned on lower noise (blue branch) to achieve self-representation alignment. We use a stop-gradient (sg) operator on the teacher to let gradients flow only through the student, and update the teacher’s parameters with an exponential moving average (ema) of the student’s parameters.
Previous studies have demonstrated that learning good internal representation can both speed up the diffusion transformer’s training and improve the quality of its generated samples. In SRA, our insight is aligning the student’s latent feature in the earlier layer conditioned on higher noise with that in the later layer conditioned on the lower noise of the teacher to conduct self-representation alignment without requiring any external representation components. As depicted in Figure 3, the goal of our simple training framework is to let the diffusion transformer not only predict noise-invariant but also align with better visual representations from itself. This operation thereby provides a simple way to enhance the representation learning in the diffusion transformer during generative training without the need to design complex representation regulation or introduce external representation foundation models.
Formally, let $f$ be the trainable student model and $f_{\ast}$ be the teacher model. Considering input noise latent, timestep, and condition to be ${\mathbf{x}}_{t}$, $t$, and $c$. Then, we can obtain the student encoder latent output ${\mathbf{y}}=f^{m}({\mathbf{x}}_{t},t,c)\in\mathbb{R}^{B\times N\times D}$, where $B,N,D>0$ are the batchsize, number of patches and the embedding dimension for $f$, and $m$ denote the output from the $m^{th}$ layer in $f$. Similarly, the output of the teacher can be expressed as ${\mathbf{y}}_{\ast}=f_{\ast}^{n}({\mathbf{x}}_{t-k},t-k,c)\in\mathbb{R}^{B\times N\times D}$. In our SRA, we set $m\leq n$, $k\geq 0$, and $0\leq(t-k)<t_{max}$or SiTs, $t_{max}=1$ and for DiTs, $t_{max}=1000$. In practice, we truncate ($t-k$) to 0 if it is less than 0.. Hence, we can conduct self-representation alignment using the teacher’s output ${\mathbf{y}}_{\ast}$ and the student’s output transform $j_{\psi}({\mathbf{y}})\in\mathbb{R}^{B\times N\times D}$, where $j_{\psi}({\mathbf{y}})$ is a projection of the student encoder output ${\mathbf{y}}$ that through a lightweight trainable MLPs head $j_{\psi}$. Notably, this projection head can be discarded optionally after training, which enables SRA to provide guidance without altering any architecture in diffusion transformers.
In particular, SRA attains self-alignment by minimizing the patch-wise distance between the teacher’s output (${\mathbf{y}}_{\ast}$) and the student’s output variant ($j_{\psi}({\mathbf{y}})$):
$\mathcal{L}_{\text{sa}}(\zeta_{s},\psi)=\mathbb{E}_{\bm{{\mathbf{x}}}_{t},t,c}\Big{[}\frac{1}{N}\sum_{i=1}^{N}\mathrm{dist}({\mathbf{y}}_{\ast}^{[i]},j_{\psi}({\mathbf{y}}^{[i]}))\Big{]},$ (5)
where $[i]$ is a patch index, $\mathrm{dist}(\cdot,\cdot)$ is a pre-defined distance calculation function, and $\zeta_{s}$, $\psi$ is the parameters of student diffusion transformer and the projection head. This objective is similar to that mentioned in REPA [81], except that we’re aiming to align with the diffusion transformer itself.
Finally, we add the abovementioned two objectives for joint learning:
$\mathcal{L}=\mathcal{L}_{\text{gen}}+\lambda\mathcal{L}_{\text{sa}},$ (6)
where $\lambda>0$ is a hyperparameter that controls the trade-off between the generation object and the self-representation alignment object.
2.3 EMA Teacher Network
In SRA, we do not need an off-the-shelf teacher to give the prior guidance. Meanwhile, using the output of the same model as the target to conduct supervision would cause training instability [12, 28]. Thus, we follow some works in self-supervised learning (SSL) [85, 8, 30] that build the teacher from past iterations of the student network using an exponential moving average (EMA) on the student weights. In specific, the updated role of EMA is $\zeta_{t}=\alpha\zeta_{t}+(1-\alpha)\zeta_{s}$, where $\alpha\in[0,1)$ is the momentum coefficient. We study different values of $\alpha$ in Section 3.2 and show that $\alpha=0.9999$ unchanged works surprisingly well in our framework. Moreover, we do not use other operations like clustering constraints [6, 7], batch normalizations [28, 67], and centering [8, 85] in SSL because we find that the training progress is already stable enough without applying these tricks.
3 Experiment
In this section, we seek to validate the effect of SRA through extensive experiments. In a specific, we mainly focus on answering the following questions:
  • How each design choice and component in SRA influence the performance? (Table 3.2)
  • Dose SRA work on different baselines across different model sizes? (Figure 4)
  • Can SRA show comparable or superior performance against other methods that leverage either the representation training paradigm or the representation foundation model? (Table 3.3)
  • Does SRA genuinely enhance the representation capacity of the baseline model, and is the generation capability indeed strongly correlated with the representation guidance? (Figure 6)
3.1 Experimental Setup
Implementation details. Unless otherwise specified, the training details strictly follow the setup in DiT [62] and SiT [56], including AdamW [54] with a constant learning rate of 1e-4, no weight decay, batchsize of 256, using the Stable Diffusion VAE [68] to extract the latent, and etc. We use ImageNet [18], where each image is preprocessed to the resolution of 256$\times$256 (denoted as ImageNet 256$\times$256), and follow ADM [19] for other data preprocessing protocols. For model configurations, we use the B/2, L/2, and XL/2 architectures introduced in the DiT and SiT papers, which process inputs with a patch size of 2. Additional experimental details, hyperparameter settings, linear probing details, and computing resources, are provided in Appendix 3.
Evaluation. We report Fréchet inception distance (FID [31]), sFID [59], inception score (IS [70]), precision (Pre.) and recall (Rec.) [44]. To ensure a fair comparison with previous methods, we also use the ADM’s TensorFlow evaluation suite  [19]with 50K samples and the same reference statistics. A detailed breakdown of each evaluation metric is included in Appendix D.
Sampler. For DiT [62], we use the DDPM sampler and set the number of function evaluations (NFE) as 250 by default. For SiT [56], we use the SDE Euler-Maruyama sampler (for SDE with $w_{t}=\sigma_{t}$) and set the NEF as 250 by default. The settings also align with those used in DiT and SiT.
Methods for Comparison. To compare performance, we utilize recent advanced methods based on diffusion models. In particular, We focus on three distinct categories of approaches: (a) Pixel diffusion: ADM [19], VDM+++++ + [40], Simple diffusion [35], CDM [33], (b) Latent diffusion with U-Net: LDM [69], and (c) Latent diffusion with transformers: DiT [62], SiT [56], SD-DiT [86], MaskDiT [84], TREAD [42], REPA [81], and MAETok [9]. Here in the diffusion transformers family, we choose to compare with DiT/SiT and their modifications with representation components involved but do not compare with works aiming at designing advanced architecture like lightningDiT [80] and DDT [74]. Detailed descriptions of each method can be found in Appendix E.
3.2 Component-Wise Analysis
Below, we provide a detailed analysis of the impact of each component. To save time, we use SiT-B/2 and train with SRA for 400K iterations for evaluation. Results are shown in Table 3.2.
Table 1: Component-wise analysis on ImageNet 256$\times$256 without classifier-free guidance (CFG). $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively. $m\xrightarrow{}n$ denotes aligning features from $m^{th}$ layer of the student with that from $n^{th}$ layer of the teacher. $[0,k_{max})$ denotes time interval $k$ is chosen randomly from 0 to $k_{max}$. PH. denotes whether to use the projection head.
[…]
SiT-B/2 Baseline [56] 33.02 43.71
$106\xrightarrow{}$ 10 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 34.85 40.58
$84\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 30.00 47.78
$104\xrightarrow{}$ 10 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 30.65 46.69
[…]
$62\xrightarrow{}$ 6 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 32.14 46.36
$82\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.31 50.13
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
$33\xrightarrow{}$ 3 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 37.08 41.54
$83\xrightarrow{}$ 8 0.0 smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 31.07 47.32
$83\xrightarrow{}$ 8 0.1 smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.55 49.01
$83\xrightarrow{}$ 8 0.2 smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 30.70 47.72
$83\xrightarrow{}$ 8 0$[0,0.1)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.38 49.32
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
$83\xrightarrow{}$ 8 0$[0,0.3)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.15 50.01
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
$83\xrightarrow{}$ 8 0$[0,0.2)$ $\ell_{1}$ Gen’s mom. $\checkmark$ 29.06 50.08
$83\xrightarrow{}$ 8 0$[0,0.2)$ $\ell_{2}$ Gen’s mom. $\checkmark$ 29.54 49.25
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Student copy $\checkmark$ 35.71 42.18
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ SSL’s mom. $\checkmark$ 33.17 44.96
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\times$ 34.23 41.07
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
Block layers for alignment. We begin by analyzing the effect of using different blocks of student and teacher for alignment. We observe that using the teacher’s last but not least few layers (e.g., 8) to regulate the student’s first layers (e.g., 3) leads to optimal performance. We assume that the first few layers need more guidance so they can catch semantically meaningful representation for subsequent generation. Meanwhile, there is a strong correlation between the quality of the representations of the teacher’s layers and the performance of the corresponding aligned student (we give the quantitative results and analysis in the latter Section 3.4). Based on these results, we set alignment layers as $83\xrightarrow{}$ 8, $166\xrightarrow{}$ 16, and $208\xrightarrow{}$ 20 for B, L, and XL models respectively as defaultFor DiTs, we set alignment layers as $73\xrightarrow{}$ 7, $146\xrightarrow{}$ 14, and $168\xrightarrow{}$ 16 by default because DiT’s discriminative behavior across layers is slightly different from SiT’s (see Figure 2 and Figure 7).
Time interval for alignment. We then study the time interval (meaning the same as $k$ in Section 2.2) used for alignment. Here, we study fixed and dynamic intervals. Notably, we find that using the teacher’s features input with lower noise level than the student’s leads to performance enhancement, and an interval value of 0.1 or the mean of 0.1 is optimal. We hypothesize that this is because lower noise levels can offer better representation guidance, but an excessively large time interval can hinder the model’s learning process, causing it to focus only on optimizing alignment loss at the expense of neglecting the generative aspects. As dynamic interval shows slightly better performance, we apply time interval as $0\sim 0.2$ in our feature experiments§§§For DiTs, we set time interval as $\lfloor[0,200)\rfloor$ by default since DiTs adopt the linear variance schedule with $t$ where $t\in\{0,1,2,...,999\}\cap\mathbb{Z}$..
Objective for alignment. We next compare three simple regression training objectives for alignment, including smooth-$\ell_{1}$, $\ell_{1}$, and $\ell_{2}$. Empirically, we find that all three objectives can lead to good performance and are stable during training. We adopt smooth-$\ell_{1}$ by default in future experiments.
Teacher network for alignment. In other generative learning studies, the EMA model is often employed solely for evaluation. However, as we need it to provide guidance during training, we study different updating methods. Here, we investigate several different strategies to build the teacher. First, we find that using teacher copied from a student would impair the performance, which testifies to our argument in Section 1. Next, we consider using the strategy used in self-supervised learning works [8, 85, 28] that updates the momentum coefficient from 0.996 to 1 during training. However, in our framework, this does not work well. Finally, we use momentum coefficient of 0.9999 unchanged, which is commonly used in other generative learning works [19, 62, 56] and find it is also suitable for our framework. Thus, we set the momentum coefficient as 0.9999 in future experiments.
Effect of projection head for alignment. We finally examine the effect of the projection head for alignment. Surprisingly, We observe that using this simple head to post-possess the student’s output is much better than directly using it to align. We hypothesize this slight operation enables the model to learn more effective hidden representations for subsequent projection head to conduct transformation for final alignment, rather than explicitly aligning the entire latent feature that could potentially disrupt the original generation field that each layer and timestep is responsible for [83, 82, 25]. Hence, we keep using the projection head in future experiments.
3.3 System-Level Comparison
In this section, a system-level comparative study is performed to assess recent diffusion-based approaches against diffusion transformers with SRA.
First, we compare the FID values between vanilla DiT or SiT and the same models trained with SRA. As shown in Figure. 4, diffusion transformers trained with SRA demonstrate substantial improvements in performance at each training step across different types , as well as various sizes. Moreover, similar to the observation in some SSL works [61, 23], we notice that the effect of SRA in a larger size model is more significant, which is probably because the larger model tends to provide richer guidance. It is also worth noting that the benefits of SRA do not saturate even when the models have already achieved a low FID score. We assume that this is likely due to the teacher’s constantly improving capacity throughout training, which allows it to provide better and better representation guidance for the student when the training goes on.
Figure 4: FID comparisons with vanilla DiTs and SiTs across different model sizes on ImageNet 256$\times$256 without classifier-free guidance (CFG).
Table 2: System-level comparison on ImageNet 256$\times$256 with Classifier-free Guidance (CFG). The best and second-best results on each metric are highlighted in bold and underlined.
[…]
Pixel diffusion
  00ADM-U 400 - 03.94 6.14 186.7 0.82 0.52
  00VDM+++++ + 560 - 02.40 - 225.3 - -
  00Simple diffusion 800 - 02.77 - 211.8 - -
  00CDM 02160 - 04.88 - 158.7 - -
Latent diffusion, U-Net
  00LDM-4 0200 LDM-VAE 03.60 - 247.7 0.87 0.48
Latent diffusion, Transformer
  00DiT-XL/2 01400 SD-VAE 02.27 4.60 278.2 0.83 0.57
  00SiT-XL/2 01400 SD-VAE 02.06 4.50 270.3 0.82 0.59
  00SD-DiT 480 SD-VAE 03.23 - - - -
  00MaskDiT 01600 SD-VAE 02.28 5.67 276.6 0.80 0.61
  00DiT + TREAD 740 SD-VAE 01.69 4.73 292.7 0.81 0.63
  00SiT + REPA 800 SD-VAE 01.42 4.70 305.7 0.80 0.65
  00SiT + MAETok 800 MAE-Tok 01.67 - 311.2 - -
  00SiT + SRA (ours) 400 SD-VAE 01.85 4.50 297.2 0.82 0.61
  00SiT + SRA (ours) 800 SD-VAE 01.58 4.65 311.4 0.80 0.63
Figure 5: Selected samples on ImageNet 256$\times$256 from the SiT-XL + SRA. We use classifier-free guidance with $w$ = 4.0. More uncurated samples are provided in Appendix J.
Finally, we provide a quantitative comparison between SiT-XL with SRA and other recent diffusion model methods using classifier-free guidance [34]. Our method already outperforms the original SiT-XL model with 1000 fewer epochs, and it is further improved with longer training. At 800 epochs, SiT-XL with SRA achieves FID of 1.58 and IS of 311.4. It is worth noting that this result is far superior to methods (e.g., MaskDiT [84]) that rely on auxiliary and complex representation training frameworks and is comparable with methods (e.g., REPA [81]) that depend heavily on an external pre-trained representation model. Moreover, because of the progressively higher-quality guidance teacher provide throughout training, we find that our method is much less likely to encounter saturation compared with REPA (we provide detailed results and analysis in Appendix F).
3.4 Ablation Study
Since SRA introduces representation guidance in an implicit way, we thus aim at testing whether representation truly matters in SRA. We give the answer with the following experiments.
Enhanced representation capacity with SRA. We first compare the representation capacity of vanilla SiT and SiT trained with SRA. As shown in Figure 6a and Figure 6b, SRA consistently improve the quality of latent representation in the diffusion transformer, as indicated by better linear probing results across different blocks and timesteps.
Tight coupling between generation quality and representation guidance in SRA. We then investigate the correlation between generation performance and the representation guidance (detailed experimental setup can be found in Appendix G) in SRA. Figure 6c reveals a strong correlation between linear probing accuracy and FID scores as the teacher network layers for alignment are varied. This finding underscores that the model’s generative capabilities are indeed closely tied to the effectiveness of the self-representation guidance mechanism.
Figure 6: We empirically investigate the effect of representations in SRA. Left and Middle: Linear probing result of vanilla SiT-XL trained for 1400 epochs and SiT-XL + SRA trained for 800 epochs. Right: Linear probing vs. FID plot of SiT-XL + SRA with different teacher’s output layers for alignment (similar plot of DiT + SRA is provided in Appendix H).
4 Related Work
Here, we highlight key related studies and defer a discussion of other relevant studies to Appendix I.
Representations guidance for diffusion transformer’s pre-training. Many recent works have attempted to introduce representation guidance in diffusion transformer training. MaskDiT [84] and SD-DiT [86] add MAE’s [29] and IBOT’s [85] training paradigm into the original DiT’s training progress. TREAD [42] design a token routing strategy with MAE loss to speed up the diffusion model’s training. REPA [81] utilizes a large-scale data pre-trained representation model for regulating the diffusion model’s latent feature. VA-VAE [80] and MAETok [9] align the latent distribution of the tokenizer with the external representation foundation model and show this alignment is beneficial to resulting diffusion models. Different from this work, we aim to look for representation guidance in the diffusion model itself and its own training paradigm.
EMA model as teacher. Unlike traditional knowledge distillation [27] that uses a well pre-trained model as the teacher, using EMA model [36] as the teacher can be seen as self-distillation because the weight of the teacher model is obtained by the weighted moving average of the student. This method often needs a feasible pretext task to succeed. For example, MoCo [30, 14] sets the teacher’s output as a momentum queue and uses contrastive learning to guide the student model; DINO [8, 61] feeds two views of images to the teacher, and the trainable student then forces the student’s output distribution to be close to that of the teacher. Our work also shares some similarities, where we set aligning the student model’s latent feature in the earlier layer conditioned on higher noise with that in the later layer conditioned on lower noise of the teacher as our pretext task for training.
5 Conclusion and Discussion
In this study, we show that diffusion transformers can provide representation guidance by themselves to boost generation performance with our proposed SRA, which aligns their latent representation in the earlier layer conditioned on higher noise to that in the later layer conditioned on lower noise to progressively enhance the representation learning without external components.
Meanwhile, a few open questions are worth discussing. First, we have observed that the effect of SRA becomes more pronounced when scaling model size. However, owing to the limitation of computational resources, we are unable to further scale the model size and data. Investigating the scalability of SRA in more complex and resource-intensive scenarios (e.g., text-to-image) will be an exciting future direction. Next, similar to other related works [81, 84, 9, 86], our method is also experiment-driven. Exploring theoretical insights into why learning a good representation is beneficial to generation will also be an exciting future direction.
Considering the simplicity and effectiveness of SRA, we believe it will facilitate more future researches that focus on unveiling the intertwined nature of visual representation and generation.
6 Acknowledgment
Our SRA manuscript referred to some of the color schemes and structural options provided in the REPA [81] manuscript. We sincerely thank REPA’s authors for their fully open-source project. We also thank Sihyun Yu, Sizhe Dang, and Zanyi Wang for the helpful discussions and suggestions during the progress of SRA project.
Appendix
Appendix A Investigation of Representations in DiT
We also perform a similar analysis with DiT like those have done in Figure 2a (PCA visualization) and Figure 2b (linear probing), the results are showed in Figure 7. In short, we also observe that the representations in DiT basically lead a process from coarse to fine when increasing block layers and decreasing noise level as SiT’s.
Figure 7: We also empirically investigate the representations in diffusion transformers across different blocks and timesteps with the original DiT-XL/2 checkpoint trained for 7M iterations. Similar to SiT, the latent representations in DiT basically follow the coarse to fine process, as block layers increase and noise level decreased.
Appendix B Descriptions for Two Types of Baseline Models
In this paper, we use DiT and SiT as our baseline models. We now provide an overview of two types of generative models that are variants of denoising autoencoders and are used to learn the target distribution. Specifically, we discuss Denoising Diffusion Probabilistic Models (DDPMs) like DiT in Section B.1 and Stochastic interpolant models like SiT in Section B.2.
B.1 Denoising Diffusion Probabilistic Models (DiT)
Denoise-based models [32, 60] aim to model the target distribution ( ${\mathbf{x}}\sim p({\mathbf{x}})$ ) by learning a gradual denoising process that transforms a Gaussian distribution ( $\mathcal{N}(\mathbf{0},\mathbf{I})$ ) into $p({\mathbf{x}})$. Formally, diffusion models learn a reverse process ( $p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ ) corresponding to a pre-defined forward process ( $q({\mathbf{x}}_{t}|{\mathbf{x}}_{0})$ ), which incrementally adds Gaussian noise to the data starting from $p({\mathbf{x}})$ over a sequence of time steps ( $t\in{1,\dots,T}$ ), with $T>0$ fixed.
For a given ${\mathbf{x}}_{0}\sim p({\mathbf{x}}_{0})$, the forward process ( $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})$ ) is defined as: $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})=\mathcal{N}({\mathbf{x}}_{t};\sqrt{1-\beta_{t}}{\mathbf{x}}_{0},\beta_{t}^{2}\mathbf{I})$, where $\beta_{t}\in(0,1)$ are pre-defined small hyperparameters. DDPM [32] formalizes the reverse process $p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ as:
$p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})=\mathcal{N}\Big{(}{\mathbf{x}}_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\big{(}{\mathbf{x}}_{t}-\frac{\sigma_{t}^{2}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\bm{\zeta}}({\mathbf{x}}_{t},t)\big{)},\mathbf{\Sigma}_{\bm{\zeta}}({\mathbf{x}}_{t},t)\Big{)},$ (7)
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}t=\prod{i=1}^{t}\alpha_{i}$, and $\bm{\epsilon}_{\bm{\zeta}}(\bm{x_{t}},t)$ is parameterized by a neural network.
The model is trained using a simple denoising autoencoder objective:
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{{\mathbf{x}}_{t},\bm{\epsilon},t}\Big{[}||\bm{\epsilon}-\bm{\epsilon}_{\bm{\zeta}}({\mathbf{x}}_{t},t)||_{2}^{2}\Big{]}.$ (8)
For the covariance term ( $\mathbf{\Sigma}_{\bm{\zeta}}({\mathbf{x}}_{t},t)$ ), DDPM[32] demonstrated that setting it as ( $\sigma_{t}^{2}\mathbf{I}$ ) with ( $\beta_{t}=\sigma_{t}^{2}$ ) is sufficient. Subsequently, Improved-DDPM[60] showed that performance can be enhanced by jointly learning ( $\mathbf{\Sigma}_{\bm{\zeta}}({\mathbf{x}}_{t},t)$ ) along with ( $\bm{\epsilon}_{\bm{\zeta}}({\mathbf{x}}_{t},t)$ ) in a dimension-wise manner using the following objective:
$\mathcal{L}_{\text{vlb}}=\exp(v\log\beta_{t}+(1-v)\log\tilde{\beta}_{t}),$ (9)
where $v$ is a component per model output dimension, and $\tilde{\beta}t=\frac{1-\bar{\alpha}{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$.
With a sufficiently large $T$ and an appropriate scheduling of $\beta_{t}$, the distribution $p({\mathbf{x}}_{T})$ approaches an isotropic Gaussian distribution. Thus, sampling is achieved by starting from random Gaussian noise and iteratively applying the reverse process ( $p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ ) to recover a data sample ${\mathbf{x}}_{0}$ [32].
B.2 Stochastic Interpolants Models (SiT)
Unlike DDPMs, flow-based models [51, 49] describe a continuous time-dependent process involving data ( ${\mathbf{x}}_{\ast}\sim p({\mathbf{x}})$ ) and Gaussian noise ( ${\mathbf{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ ) over $t\in[0,1]$:
${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{0}+\sigma_{t}\bm{\epsilon},\quad\alpha_{0}=\sigma_{1}=1,\,\,\alpha_{1}=\sigma_{0}=0,$ (10)
with $\alpha_{t}$ and $\sigma_{t}$ being decreasing and increasing functions of $t$, respectively. The process is governed by a probability flow ordinary differential equation (PF ODE):
$\dot{{\mathbf{x}}}_{t}={\mathbf{v}}({\mathbf{x}}_{t},t),$ (11)
where the distribution of the ODE at time $t$ matches the marginal distribution $p_{t}({\mathbf{x}})$.
and can be approximated by a model ${\mathbf{v}}_{\bm{\zeta}}({\mathbf{x}}_{t},t)$ trained to minimize the objective:
$\mathcal{L}_{\text{velocity}}=\mathbb{E}_{{\mathbf{x}}_{0},\bm{\epsilon},t}\Big{[}||{\mathbf{v}}_{\bm{\zeta}}({\mathbf{x}}_{t},t)-\dot{\alpha}_{t}{\mathbf{x}}_{0}-\dot{\sigma}_{t}\bm{\epsilon}||^{2}\Big{]}.$ (12)
This also corresponds to a reverse stochastic differential equation (SDE) given by:
$d{\mathbf{x}}_{t}={\mathbf{v}}({\mathbf{x}}_{t},t)dt-\frac{1}{2}w_{t}{\mathbf{s}}({\mathbf{x}}_{t},t)dt+\sqrt{w_{t}}d\bar{\mathbf{w}}_{t},$ (13)
As shown in [2], any functions $\alpha_{t}$ and $\sigma_{t}$ that satisfy the following three conditions:
$1.~{}$ $\alpha_{t}^{2}+\sigma_{t}^{2}>0,\,\,\forall t\in[0,1]$
$2.~{}$ $\alpha_{t}~{}\text{and}~{}\sigma_{t}~{}\text{are differentiable},\,\,\forall t\in[0,1]$
$3.~{}$ $\alpha_{1}=\sigma_{0}=0,~{}\alpha_{0}=\sigma_{1}=1,$
lead to an unbiased interpolation process between ${\mathbf{x}}_{0}$ and ${\mathbf{\epsilon}}$. Example choices include linear interpolants ($\alpha_{t}=1-t$, $\sigma_{t}=t$) or variance-preserving (VP) interpolants ($\alpha_{t}=\cos(\frac{\pi}{2}t)$, $\sigma_{t}=\sin(\frac{\pi}{2}t)$) [56].
An advantage of stochastic interpolants is that the diffusion coefficient ( $w_{t}$ ) can be independently selected during sampling with the reverse SDE, even after training. This capability simplifies the design space compared to score-based diffusion models [39].
Appendix C Hyperparameter and More Implementation Details
Table 3: Defualt hyperparameter setup. Unless other otherwise specified, we use these sets of hyperparameters for different models. In our component-wise analysis experiment, settings are also kept the same except those we point out in Table 3.2.
SiT-B SiT-L SiT-XL DiT-B DiT-L DiT-XL
Architecture
Input dim. 32$\times$32$\times$4 32$\times$32$\times$4 32$\times$32$\times$4 32$\times$32$\times$4 32$\times$32$\times$4 32$\times$32$\times$4
Patch size 2 2 2 2 2 2
Num. layers 12 24 28 12 24 28
Hidden dim. 768 1024 1152 768 1024 1152
Num. heads 12 16 16 12 16 16
SRA
Alignment blocks $83\xrightarrow{}$ 8 $166\xrightarrow{}$ 16 $208\xrightarrow{}$ 20 $73\xrightarrow{}$ 7 $146\xrightarrow{}$ 14 $168\xrightarrow{}$ 16
Alignment time interval $[0,0.2)$ $[0,0.2)$ $[0,0.2)$ $\lfloor[0,200)\rfloor$ $\lfloor[0,200)\rfloor$ $\lfloor[0,200)\rfloor$
Objective smooth-$\ell_{1}$ smooth-$\ell_{1}$ smooth-$\ell_{1}$ smooth-$\ell_{1}$ smooth-$\ell_{1}$ smooth-$\ell_{1}$
EMA decay 0.999 0.999 0.999 0.999 0.999 0.999
Using projection head $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Optimization
Batch size 256 256 256 256 256 256
Optimizer AdamW AdamW AdamW AdamW AdamW AdamW
lr 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
$(\beta_{1},\beta_{2})$ (0.9, 0.999) (0.9, 0.999) (0.9, 0.999) (0.9, 0.999) (0.9, 0.999) (0.9, 0.999)
Interpolants or Denoising
$\alpha_{t}$ $1-t$ $1-t$ $1-t$
$\sigma_{t}$ $t$ $t$ $t$
$w_{t}$ $\sigma_{t}$ $\sigma_{t}$ $\sigma_{t}$
T - - - 1000 1000 1000
Training objective v-prediction v-prediction v-prediction noise-prediction noise-prediction noise-prediction
Sampler Euler-Maruyama Euler-Maruyama Euler-Maruyama DDPM DDPM DDPM
Sampling steps 250 250 250 250 250 250
Guidance Scale - - 1.8 (if used) - - -
More implementation details. We implement our models based on the original DiT and SiT implementation. To speed up training and save GPU memory, we use mixed-precision (fp16) with a gradient clipping and FusedAttention [17] operation for attention computation. We also pre-compute compressed latent vectors from raw pixels via stable diffusion VAE [68] and use these latent vectors. We do not apply any data augmentation, but we find this does not lead to a big difference, as similarly observed in MaskDiT [84] and REPA [81]. We also use stabilityai/sd-vae-ft-ema for encoding images to latent vectors and decoding latent vectors to images. For the projection head used for nonlinear transformation, we use two-layer MLP with SiLU activations [21]. The $\lambda$ is set to 0.04 for DiT and 0.2 for SiT initially and followed the rules: $\lambda=\lambda_{\text{init}}*0.1^{\left(\frac{N_{e}-150}{1000}\right)},$ where $N_{e}>150$ is the number of epoch, to change after 150 epochs for ensuring $\mathcal{L}_{\text{gen}}$ and $\mathcal{L}_{\text{sa}}$ at the same scale.
When using SiT-XL to generate images with classifier-free guidance [34], the guidance interval introduced in the study [43] with the same setting used in REPA [81] is applied, which has been demonstrated to yield a slight performance improvement.
Linear probing. We follow the setup used in REPA [81] and I-DAE [13], and use the code-base of ConvNeXt [53] to conduct the experiment. Specifically, we use AdaptiveAvgPooling and a batch normalization layer to process the output of the latent feature by the model, then train a linear layer for 80 epochs. The batch size is set to 4096 with a cosine decay learning rate scheduler, where the initial learning rate is set to 0.001.
Computing resources. We use 8 NVIDIA A100 80GB GPUs or 8 NVIDIA L40S 48GB GPUs for training largest model (XL); and use 4 NVIDIA A100 80GB GPUs or 4 NVIDIA L40S 48GB GPUs for training smaller model (L, B), our training speed is about 2.12 step/s with a global batch size of 256. When sampling, we use either NVIDIA A100 80GB GPUs or NVIDIA L40S 48GB GPUs or NVIDIA RTX 4090 24GB GPUs to obtain the samples for evaluation.
Appendix D Evaluation Metric
In this section, we define the key metrics used to assess the performance of our model. Each metric is summarized below:
  • Fréchet Inception Distance (FID) [31]
    Purpose: Measures the similarity between the feature distributions of real and generated images.
    Methodology: Uses the Inception-v3 network [73] to extract features. Assumes both feature distributions are multivariate Gaussian and computes the Fréchet distance between them.
    Interpretation: Lower FID scores indicate better similarity (and thus higher-quality generated images).
  • Spatial Fréchet Inception Distance (sFID) [59]
    Purpose: Extends FID by incorporating spatial information to better capture the structural fidelity of generated images.
    Methodology: Computes FID using intermediate spatial features (rather than global features) from the Inception-v3 network.
  • Inception Score (IS) [70]
    Purpose: Evaluates the quality and diversity of generated images.
    Methodology: Uses the Inception-v3 network to compute class probabilities (logits) for generated images. Measures the KL-divergence between the marginal class distribution of generated images and the conditional class distribution of a single image (after softmax normalization).
    Interpretation: Higher IS values indicate both high image quality (confident predictions) and diversity (uniform marginal distribution).
  • Precision and Recall for Distributions (Precision/Recall) [44]
    Purpose: Evaluates the trade-off between sample quality (precision) and distribution coverage (recall).
    Methodology: Precision: Fraction of generated images deemed realistic by a classifier (relative to real images). Recall: Fraction of the real image manifold covered by generated samples.
Appendix E Methods for Comparison
Next, we explain the key ideas behind the methods used for evaluation and comparison.
  • ADM [19] enhances U-Net-based architectures for diffusion models and introduces classifier-guided sampling, a technique used to balance the quality-diversity tradeoff and improve overall performance.
  • VDM++ [40] improves diffusion model training efficiency through an adaptive noise scheduling mechanism that dynamically adjusts noise levels during optimization.
  • Simple Diffusion [35] targets high-resolution synthesis by simplifying both network architectures and noise schedules through systematic exploration of lightweight design choices.
  • CDM [33] achieves high-fidelity generation via a cascaded pipeline: training low-resolution diffusion models first, then progressively refining details through super-resolution diffusion stages.
  • LDM [69] accelerates training while maintaining generation quality by operating in compressed latent spaces, modeling image distributions at reduced dimensionality compared to pixel space.
  • DiT [62] employs a pure transformer backbone for diffusion models, incorporating AdaIN-zero modules as an important component modification against vanilla vision transformer [20] of its architecture.
  • SiT [56] provides an extensive analysis of how the training efficiency of DiTs can be improved by transitioning from discrete diffusion to continuous flow-based modeling, which can be seen as a flow-based version of DiT.
  • SD-DiT [86] leverages IBOT’s [85] training paradigm that combines DINO loss [8] and BEIT loss [4] for efficiently training diffusion transformers.
  • MaskDiT [84] proposes an asymmetric encoder-decoder scheme for efficient training of diffusion transformers, where they train the model with an auxiliary mask reconstruction task similar to MAE [29].
  • TREAD [42] introduces a dynamic token routing strategy combined with the mask reconstruction task similar to MAE [29] and MaskDiT [84] to accelerate the training of diffusion models.
  • REPA [81] achieves significant improvements in both training efficiency and generation quality by aligning the latent feature of the diffusion model with that of a large-scale data pre-trained representation model (e.g., DINOv2 [61]).
  • MAETok changes the SD-VAE to MAE-Tok, which is trained with auxiliary mask reconstruction loss and aligns loss with three representation targets (HOG’s [15], DINOv2’s [61], and CLIP’s [65]) and obtains diffusion transformer with better generation performance.
Appendix F Convergence Speed Against REPA
Here we provide a quantitative comparison of convergence speed with SiT-XL baseline and SiT-XL + REPA [81]. It is worth noting that REPA leverages a powerful representation foundation model whose training data and resources are far beyond training a diffusion transformer on ImageNet.
As shown in Figure 8, although REPA converges quickly at the beginning of training due to its large-scale, pre-trained representation model, the performance saturates after about 200 epochs. On the contrary, in SRA, since the teacher’s ability is constantly enhanced during training and thus better and better guidance is provided, continuous performance improvement can be achieved. From another perspective, the representation guidance signal in REPA is fixed in every timestep and epoch, while our supervision signal is dynamically changing. This provides a wider variety of learning cues to help the model continuously improve its performance.
Figure 8: Training epochs vs. FID plot without classifier-free guidance (CFG). The benefit of REPA encounters saturation after approximately 200 epochs, while our method provides continuous performance improvement.
Appendix G Detailed Setup of Ablation Study
We now give the detailed experimental setup of Figure 6c in our ablation study. The accuracy rate of the horizontal axis in the figure is obtained by using the linear probing results of the original SiT-XL/2 checkpoint training for 7M iterations, while the vertical axis is the FID evaluation result of training 400K iterations with SRA without classifier-free guidance (CFG). Since we do not introduce any representation component and use the representation supervision signal only in the generative training process, we consider this experiment to validate the effectiveness of our approach.
Appendix H Ablation Results of DiT
We also perform a similar analysis with DiT like those have done in Figure 6c, the results are showed in Figure 9. In short, we also observe that the generative capability of DiT with SRA is indeed strongly correlated with the representation guidance as observed in SiT with SRA.
Figure 9: We also investigate the correlation between generation performance and the representation guidance of DiT + SRA. A similar tight coupling can also be seen.
Appendix I More Discussion on Related Work
We now provide a detailed literature review of other related work.
Diffusion transformers. Currently, the diffusion model is progressively transitioning from a U-Net-based architecture to a Transformer-based one, owing to the superior scalability of the latter. At the beginning, U-ViT [3] shows transformer-based backbones with skip connections can be an effective backbone for training diffusion models. Then, DiT [62] shows skip connections are not even necessary components, and a pure transformer architecture can be a scalable architecture for training denoise-based models. Based on DiT, SiT [56] shows the model can be further improved with continuous stochastic interpolants [2]. Moreover, Stable diffusion 3 [22], Lumina-Image 2.0 [64] and FLUX 1 [45] show pure transformers can be scaled up for challenging text-to-image generation, and this characteristic is also verified by Sora [5], CogvideoX [79] and Wan [76] in text-to-video field. Our work focuses on improving the training of DiT (and SiT) architecture based on our proposed self-representation alignment technique.
Exploring representation capacity of diffusion models. With the success of diffusion models to generate detailed images, many works have attempted to test whether discriminative semantic information can be found in diffusion models. GD [58] and DDAE [77] first observe that the intermediate representations of diffusion models have discriminative properties. Driving from this finding, I-DAE [13] deconstructs diffusion models to be a self-supervised Learner. Moreover, Repfusion [78], DiffusionDet [11], and DreamTeacher [46] use diffusion models to perform various downstream tasks (e.g., semantic segmentation and object detection). Our work also tries to explore the representation capacity of the diffusion model, but we focus on leveraging the representations in diffusion transformers to enhance their generation capacity.
Applying representation guidance to other generation tasks. In addition to pre-training of the class-conditional diffusion model, applying representation guidance can benefit other generation tasks. For example, lbGen [37] utilizes text features from CLIP [65] as a low-biased reference to regulate diffusion model for low-biased dataset synthetics. RCG [47] focuses on unconditional generation, which first use the features from a self-supervised image encoder to train a representation generator and subsequently used it output as the ‘label’ for image generation by a second generator. Diff-AE [63] uses a learnable encoder for discovering the high-level semantics and a diffusion model for modeling stochastic; this dual-encoding improves the realism of the generated images on attribute manipulation and image interpolation tasks. Different from these works, our study focuses on the class-conditional diffusion transformer’s pre-training and exploiting representation guidance in itself and its own training paradigm. But we also believe our plug-and-play method can be easily applied to other tasks with benefits.
Knowledge distillation for diffusion model. Knowledge distillation [27] is also widely used in diffusion models. Its purposes can roughly be divided into two types: improving generation performance and accelerating inference sampling. To achieve the first goal, A more powerful pretrained model is often used to guide the diffusion model during training [16, 81, 24]. For example, TinyFusion [24] combines progressive distillation with architecture-specific optimizations for U-Net backbones, enabling deployment on edge devices. REPA [81] distills the knowledge from a large-scale, pretrained representation foundation model [61, 65] and finds that this distillation can improve both training efficiency and generation quality. Meanwhile, to achieve the second goal, the student and teacher models will be initialized by pre-trained models (here it can be one model or two different models), then the training target is to make the prediction of the student model with fewer steps conform to that of the teacher model with multiple steps [55, 66, 57, 71], thereby achieving the result of accelerated sampling and generation speed. For example, CM [71] constrains the student’s outputs of adjacent points on a sampling path to be consistent with teacher’s. LCM [55] leverages this idea in the latent space. Our study also have some similarities, while we do not need a pretrained model and improve the generation performance of diffusion transformers by applying proposed SRA.
Appendix J More Qualitative Results
Below we show some uncurated generation results on ImageNet 256$\times$256 from the SiT-XL + SRA. We use classifier-free guidance with $w$ = 4.0.
Figure 10: Uncurated samples of loggerhead turtle (class label: 33).
Figure 11: Uncurated samples of sulphur-crested cockatoo (class label: 89).
Figure 12: Uncurated samples of golden retriever (class label: 207).
Figure 13: Uncurated samples of white fox (class label: 279).
Figure 14: Uncurated samples of tiger (class label: 292).
Figure 15: Uncurated samples of red panda (class label: 387).
Figure 16: Uncurated samples of acoustic guitar (class label: 402).
Figure 17: Uncurated samples of balloon (class label: 417).
Figure 18: Uncurated samples of baseball (class label: 429).
Figure 19: Uncurated samples of fire truck (class label: 555).
Figure 20: Uncurated samples of laptop (class label: 620).
Figure 21: Uncurated samples of ice cream (class label: 928).
Figure 22: Uncurated samples of cheeseburger (class label: 933).
Figure 23: Uncurated samples of cliff drop-off (class label: 972).
Figure 24: Uncurated samples of coral reef (class label: 973).
Figure 25: Uncurated samples of lakeside (class label: 975).